OpenStudy (h0pe):

"Modulo m graph paper" consists of a grid of m^2 points, representing all pairs of integer residues (x,y) where $0\le x

OpenStudy (rational):

use ` \( some latex mess\) ` to write inline latex

OpenStudy (h0pe):

sorry I use that format for other sites let me fix it

OpenStudy (h0pe):

"Modulo m graph paper" consists of a grid of m^2 points, representing all pairs of integer residues (x,y) where \[0\le x<m\]. To graph a congruence on modulo m graph paper, we mark every point $(x,y)$ that satisfies the congruence. For example, a graph of \[y\equiv x^2\pmod 5\] would consist of the points (0,0), (1,1), (2,4), (3,4), and (4,1). The graphs of $$y\equiv 5x+2\pmod{16}$$ and $$y\equiv 11x+12\pmod{16}$$ on modulo 16 graph paper have some points in common. What is the sum of the x-coordinates of those points?

OpenStudy (rational):

so you're given a congruence system : \[y\equiv 5x+2\pmod{16}\\~\\y\equiv 11x+12\pmod{16}\] and asked to find the sum of x values of the solutions

OpenStudy (rational):

simply subtract both the congruences so that "y" disappears

OpenStudy (h0pe):

basically just 11x+12-5x+2?

OpenStudy (h0pe):

or 6x+10 (mod 16)

OpenStudy (rational):

Yes all that equals 0 in mod 16

OpenStudy (rational):

because left hand side is 0

OpenStudy (rational):

\[y\equiv 5x+2\pmod{16}\\~\\y\equiv 11x+12\pmod{16}\] subtracting you get \[0\equiv 6x+10\pmod{16}\]

OpenStudy (rational):

which is same as \[6x\equiv -10\pmod{16}\]

OpenStudy (rational):

same as \[6x\equiv 6\pmod{16}\]

OpenStudy (h0pe):

so do I now find the inverse of 6?

OpenStudy (h0pe):

mod 16

OpenStudy (rational):

there wont be an inverse for 6 in mod 16

OpenStudy (h0pe):

then what..?

OpenStudy (rational):

maybe try finding it, you will see why there is no inverse

OpenStudy (h0pe):

ok

OpenStudy (h0pe):

then how do I solve it?

OpenStudy (rational):

simply divide 6 through out recalling the congruence property : \[\large ax\equiv a \pmod{n} ~\implies ~ x \equiv 1 \pmod{\frac{n}{\gcd(n,a)}}\]

OpenStudy (rational):

\[6x\equiv 6\pmod{16}~ \implies~ x\equiv 1\pmod{8}\]

OpenStudy (rational):

so solutions to given system are \(x=8k+1\) where \(k\in \mathbb{Z}\)

OpenStudy (rational):

since you want only the solutions between \(0\le x\lt 16\), you get two x values : \[x=1, ~9\]

OpenStudy (rational):

add them up

OpenStudy (h0pe):

10, thank you!

OpenStudy (rational):

yw